OpenStudy (anonymous):

Solve the given proportion. 1/9 =x+3/2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

\[\frac{ 1 }{ 9 } = \frac{ x+3 }{ 2 }\]

OpenStudy (anonymous):

Sorry, on bad internet. Ok so cross multiply and set them equal. 9x+27=2

OpenStudy (anonymous):

o ok thanks

OpenStudy (anonymous):

Yep, Np :)